Consider the following hypothesis test.
H0: µ1 - µ2 = 0
Ha: µ1 - µ2 ≠ 0
The following results are for two independent samples taken from two populations.
Sample 1: n1 = 80, mean1 = 104, standard deviation1 = 8.4
Sample 2: n2 = 70, mean2 = 106, standard deviation2 = 7.6
What is the value of the test statistic?
What is the p-value?
With α = .05, what is your hypothesis testing conclusion?

Answer by Boreal(15235)

SE = sqrt (s1^2/n1 + s2^2/n2)
=sqrt(70.56/80+57.76/70)=sqrt(1.707)=1.31
test statistic is a t(0.975, df=148)=1.98
reject for |t|>1.98
t=diff means/SE=-2/1.31=-1.53 fail to reject
p-value is 0.13